# Dynamical scheme for computing the mass parameter of a system in a medium

### v1

@@ III.2 Quadrupole moment oscillations
Applying such potential results in oscillations between oblate and prolate shapes consecutively.
Figure 5: a) Oscillations of the quadrupole moment $Q_{20}$ as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8).
@@
where $\kappa=\rho_{\textrm{in}}/\rho_{\textrm{out}}$, with $\rho_{\textrm{out}}$ being density of neutrons far from the impurity and $\rho_{\textrm{in}}$ being density of protons and neutrons inside the impurity. We compare this prediction with our results in Fig. 6b). We find that for higher densities, the results given by the hydrodynamic method stand in good agreement with the results that we obtained. On the other hand, for densities below 0.03 $0.03\textrm{~{}fm}^{-3}$, the values obtained from hydrodynamics overestimate our predictions, in the low-density limit by a factor of about two. The discrepancies arise, e.g., from the estimation of the radius of impurity — in the hydrodynamical model, we assume a sharp radius of the impurity.
@@ Acknowledgements
Data underlying the study and detailed instructions on how to reproduce the results are available via Zenodo repository Zdanowicz et al. (2025).
@@
[…]
@@
$\bar{\rho}$ $\rho_{Bn}$ $M_{CM}$ (A=1fm) $M_{CM}$ (A=10 fm) $\gamma_{\textrm{CM}}$ (A=10 fm) $M_{Q_{20}}$ $\hbar\omega_{Q_{20}}$
[…]
0.0023 0.0016 147.17 152.09 5.05e-06 0.104 21.7
0.0058 0.0045 134.63 142.78 2.02e-05 0.107 22.2
0.0104 0.0084 161.81 171.63 6.03e-05 0.0878 17.2
0.0148 0.0120 151.49 162.52 8.74e-05 0.0870 17.9
0.0187 0.0152 170.91 186.30 7.77e-05 0.0777 16.2
0.0237 0.0193 163.46 180.44 1.38e-04 0.0756 16.1
0.0267 0.0217 162.98 184.11 1.70e-04 0.0719 15.3
0.0300 0.0244 165.54 185.64 2.05e-04 0.0675 14.5
0.0338 0.0276 160.94 178.12 2.74e-04 0.0655 15.6
0.0428 0.0351 147.19 160.73 3.37e-04 0.0567 12.3
0.0510 0.0422 147.15 163.48 5.24e-04 0.0508 11.3
Table 1: The consecutive columns describe the layer of the inner crust of density $\bar{\rho}$ with bulk neutron density $\rho_{Bn}$. We provide the effective mass $M_{\textrm{eff}}$ of a zirconium nucleus for amplitudes of oscillations equal to 1 fm and 10 fm, as well as its susceptibility for deformation $M_{Q_{20}}$. The damping coefficient for A=10fm oscillations is given by $\gamma_{\textrm{CM}}$. The quadrupole mode energy $\hbar\omega_{Q_{20}}$ is depicted in that last column.

### v2

@@ III.2 Quadrupole moment oscillations
Applying such potential results in oscillations between oblate and prolate shapes consecutively. As in the case of center-of-mass oscillations, the lack of strict periodicity in $U_{\textrm{p}}$, as required by our code, does not pose a problem, since the potential couples only to the proton density, which vanishes at the boundaries of the simulation domain.
Figure 5: a) Oscillations of the quadrupole moment $Q_{20}$ as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8), $\omega^{2}(k)=(0.1023)^{2}+\frac{k}{0.0973}$.
@@
[…]
@@ Acknowledgements
The data that support the findings of this article, with detailed instructions on how to reproduce the results, are openly available Zdanowicz et al. (2025).
@@
$T=\$$\frac{1}{2}m_{n}\rho_{\textrm{in}}\int_{r=R_{0}}\Phi_{\textrm{in}}\frac{\partial\Phi_{\textrm{in}}}{\partial r}R_{0}^{2}\,d\Omega$
$-\frac{1}{2}m_{n}\rho_{\textrm{out}}\int_{r=R_{0}}\Phi_{\textrm{out}}\frac{\partial\Phi_{\textrm{out}}}{\partial r}R_{0}^{2}\,d\Omega.$ (35)
@@
$\bar{\rho}$ $\rho_{Bn}$ $R_{0}$ $M_{\textrm{c.m.}}$ (A=1fm) $M_{\textrm{c.m.}}$ (A=10 fm) $\gamma_{\textrm{c.m.}}$ (A=10 fm) $M_{Q_{20}}$ $\hbar\omega_{Q_{20}}$
[ $\textrm{~fm}^{-3}] [$ $\textrm{~fm}^{-3}] [$ $\textrm{~fm}] [$$m_{n}] [$$m_{n}] [$$c\textrm{~fm}^{-1}] [$$\textrm{MeV}\textrm{~fm}^{-2}c^{-2}] [$$\textrm{MeV}]$
0.0023 0.0016 6.16 147.17 152.09 5.05e-06 0.104 21.7
0.0058 0.0045 5.96 134.63 142.78 2.02e-05 0.107 22.2
0.0104 0.0084 6.53 161.81 171.63 6.03e-05 0.0878 17.2
0.0148 0.0120 6.39 151.49 162.52 8.74e-05 0.0870 17.9
0.0187 0.0152 6.78 170.91 186.30 7.77e-05 0.0777 16.2
0.0237 0.0193 6.72 163.46 180.44 1.38e-04 0.0756 16.1
0.0267 0.0217 6.76 162.98 184.11 1.70e-04 0.0719 15.3
0.0300 0.0244 6.85 165.54 185.64 2.05e-04 0.0675 14.5
0.0338 0.0276 6.83 160.94 178.12 2.74e-04 0.0655 15.6
0.0428 0.0351 6.73 147.19 160.73 3.37e-04 0.0567 12.3
0.0510 0.0422 6.84 147.15 163.48 5.24e-04 0.0508 11.3
Table 1: The consecutive columns describe the layer of the inner crust of density $\bar{\rho}$ with bulk neutron density $\rho_{Bn}$. We provide the radius of the zirconium nucleus $R_{0}$, computed according to formula (40). The effective masses $M_{\textrm{c.m.}}$ are provided for runs with amplitudes of oscillations equal to 1fm and 10fm. $M_{Q_{20}}$ indicates the mass parameter related to the quadrupole deformation. The damping coefficient for A=10fm oscillations is given by $\gamma_{\textrm{c.m.}}$. The quadrupole mode energy $\hbar\omega_{Q_{20}}$ is depicted in that last column. The number of bound neutrons slightly depends on the amplitude. However, it can be easily obtained by subtracting the number of protons (Z=40) from the column $M_{\textrm{c.m.}}$.